OpenStudy (anonymous):

e^(ln2x^5) evaluate or simplify without using a calculator

OpenStudy (aum):

Is it ln(2) * x^5 or ln(2x^5) ?

OpenStudy (anonymous):

2nd one

OpenStudy (aum):

\[\Large e^{\ln(A)} = A \]

OpenStudy (aum):

\[ \Large e^{\ln(2x^5)} = 2x^5 \]